A BACTERIA CULTURE STARTED WITH A COUNT OF 720 AT 8:00 A.M., AND AFTER t HOURS IS EXPECTED TO GROW TO f(t)= 720(3/2)^t. ESTIMATE THE NUMBER OF BACTERIA IN THE CULTURE AT 11:00 A.M. THE SAME DAY.
Answer by stanbon(48535)

A BACTERIA CULTURE STARTED WITH A COUNT OF 720 AT 8:00 A.M., AND AFTER t HOURS IS EXPECTED TO GROW TO f(t)= 720(3/2)^t. ESTIMATE THE NUMBER OF BACTERIA IN THE CULTURE AT 11:00 A.M. THE SAME DAY.
f(t)=720(3/2)^t
8:00AM to 11:00AM is 3 hrs.
f(3)=720(3/2)^3
f(3)=720(27/8)
f(3)=2430 bacteria by 11:00AM
--------
Cheers,
Stan H.